OpenStudy (anonymous):

A company overvalues its shares by 25 percent. If the quoted price of a share is $300, what is the true and fair price of the share?

OpenStudy (anonymous):

The variable you're trying to solve for is the true and fair price of the share, let's call this x. We know that the company over-values their shares by 25 percent. This means they are charging 25% of the fair price above the fair price, or: x + 25% of x = $300, or x + 0.25*x = $300. By combining the x terms, you will get: 1.25*x = $300 Now just solve for x: x = $300/(1.25)